# Imaginary gap-closed points and non-Hermitian dynamics in a class of dissipative systems

### v1

@@ I Introduction
On the other hand, the NHSE also induces novel dynamical phenomena in open quantum systems. These include damping behavior and diverging relaxation time in Liouvillian dynamics [100, 101, 102, 103, 104], directional amplification of signals [105, 106], self-healing of skin modes [107], and directional invisibility of scattered wave-packet [108, 109]. Intriguingly, non-Hermitian dynamics exhibit boundary condition independence in the thermodynamic limit [110]. However, in finite-size systems, energy spectrums under open boundary conditions (OBCs) show dramatic differences from those under periodic boundary conditions (PBCs) due to the NHSE [69, 97, 98]. Thus, one may question which non-Hermitian Hamiltonians determine the dynamical evolution of the systems. Furthermore, the boundary’s role is also crucial to understanding non-Hermitian dynamics.
Motivated by the above questions, we first analyze a general non-Hermitian model whose non-Hermiticity comes from the onsite dissipation. We prove imaginary gap-closed (IGC) points are only determined by the Hermitian part of the Hamiltonian and irrelevant to the dissipation. Furthermore, IGC states are populated on non-dissipative sites and can be viewed as dark modes. When focusing on a class of one-dimensional dissipative models, the PBC spectrums exhibit IGC points, while their OBC counterparts lack such points as eigenstates feature the NHSE. By introducing the non-Hermitian quantum walk, we investigate the relation between IGC points and the scaling behavior of bulk loss probability. As a result, these IGC points induce power law in the bulk. Moreover, an associated boundary phenomenon, termed ”edge burst,” emerges. Specifically, when a quantum particle is initially prepared on a specific bulk site, there is a prominent peak in the loss probability at the edge. We discuss this phenomenon in different quantum walk scenarios and the impact of various parameters, showing the universality of the correlation between edge burst and IGC points. Finally, we analyze the Liouvillian dynamics and demonstrate that damping matrices of these dissipative chains under PBCs can have Liouvillian gapless points, implying an algebraic convergence towards the steady state in long-time dynamics.
@@ II Non-Hermitian Model and universal imaginary gapless modes
Let us first consider a general non-Hermitian lattice model [Fig. 1], whose non-Hermiticity totally comes from the onsite dissipation. We classify the lattice sites into two categories: non-dissipative sites and dissipative sites, denoted by $(h,i)$ and $(nh,i)$, respectively. Then, the Hamiltonian of this system $H$ can be divided into three parts,
@@
where $H_{h}$ is the interactions within non-dissipative sites, $H_{nh}$ is the interactions within dissipative sites, and $H_{inter}$ is the interactions between non-dissipative sites and dissipative sites,
@@
$H_{inter}=$ $\sum_{i,j}\left(c_{ij}\ket{nh,i}\bra{h,j}+h.c.\right).$
$H_{h}$ and $H_{inter}$ are Hermitian operators, while $H_{nh}$ is non-Hermitian.
@@
The wave functions of IGC states are only populated on non-dissipative sites, i.e.
@@
where $-i\gamma_{j}$ is the dissipative rate of site $(nh,j)$, and $\gamma_{j}>0$. Notice that the LHS of Eq. (6) also equals to
@@
$(H_{h}-E_{IGC})\ket{IGC}=0,\quad H_{inter}\ket{IGC}=0,$ (9)
which means that IGC states are also the eigenstates of a system composed of non-dissipative sites and satisfy the connection conditions.
In summary, IGC states can be described as dark modes characterized by the absence of a population on dissipative sites. Notably, IGC states are the eigenstates of $H_{h}$, satisfying the connection conditions $H_{inter}$, independent of $H_{nh}$.
@@ III Imaginary Gapless modes and dynamics of Quantum walk
where $k$ is a real number within the interval $\left[0,2\pi\right)$. The specific value of $k$ is determined by the connection condition obtained in Eq. (9)
@@
At $k=0$, the LHS of Eq. (12), denoted as $F(k)$, attains its maximum value $\sum_{m=0}^{n}t_{m}>0$. So if the minimum value of $F(k)$ is smaller than or equal to zero, Eq. (12) has real roots, and the energy spectrum is IGC.
@@
Notably, the system under OBC lacks IGC states. This is because any IGC state would also be an eigenstate of the Hermitian $H_{h}$, which is an extended state, contradicting the NHSE’s localized nature under OBC. Alternatively, in systems with the NSHE, OBC spectrums are enclosed by their PBC counterparts [98, 109]. For this dissipative lattice, the PBC spectrums’ maximum imaginary part is zero, rendering the OBC spectrums always imaginary-gapped. An example is shown in Fig. 3(a).
Although substantial differences exist in the spectrum and eigenstates between OBC and PBC for this model, the dynamics within the bulk exhibit similarity for both cases. This similarity is evident from the Schrödinger equation $i(d/dt)\ket{\psi(t)}=H\ket{\psi(t)}$. Under PBC, the Hamiltonian is given by $H_{PBC}=H_{OBC}+\delta H$. Here, $\delta H$ represents the coupling term that connects the two boundaries. When a wave function |$\ket{\psi(t)}$ predominantly resides in the bulk, with negligible population at the boundaries, $\delta H\ket{\psi(t)}\approx 0$, and then consequently, the Schrödinger equations for OBC and PBC become effectively equivalent in this approximation. This feature is exemplified in Fig. 3(b). In the following, we show that the dynamical evolution properties of quantum walks under OBCs can be elucidated by IGC points of the PBC spectrums.
Figure 3: The model with only the nearest A-B coupling (a) Energy spectrums with $t_{0}=0.3$. Blue: OBC; Red: PBC. (b) The bulk distribution of the relative loss probabilities $P_{x}/P_{130}$ in a double logarithmic plot with $t_{0}=0.3$ for system size $L=200$, initial position $x_{0}=150$. (c), (d) The distribution of $P_{x}$ for a walker initiated at $x_{0}=150$ with $L=200$ under OBC. $t_{0}=0.3$ for (c) and $t_{0}=0.6$ for (d). (e) Relative height $P_{\rm{edge}}/P_{\min}$ and (f) the edge loss probability $P_{\rm{edge}}$ with chain length $L=200$, and $x_{0}$ varying from 40 to 160. $t_{0}=0.3$ for (c) and $t_{0}=0.6$ for (d) in a logarithmic plot. Throughout (a)-(f), $\phi=\pi/2$, $t_{1}=0.5$, $t_{p}=0.5$, and $\gamma=0.5$ are fixed.
@@
Note that $\Sigma_{x}P_{x}=1$ is satisfied under the initial-state normalization $\left\langle\psi(0)|\psi(0)\right\rangle=1$. Intuitively, the distribution $P_{x}$ features a peak centered on the site $x_{0}$, displaying left-right asymmetry attributable to the NHSE, which causes the wave function to move leftward preferentially. However, under certain parameters, a prominent peak, named edge burst, emerges at the left edge [115]. In a recent paper, Xue et al. investigated this model with the nearest A-B hopping term and uniform dissipation, i.e., $\gamma_{x}$ is a constant, revealing that the edge burst phenomenon results from the interplay between the IGC spectrum under PBC and the NHSE [114].
@@ IV Impact of parameters
The above discussion shows that IGC points of PBC spectrums are crucial to the non-Hermitian dynamics. In this section, we explore the quantum walk across various parameters, demonstrating the universality of the correlation between edge burst and IGC points. This correlation remains independent of the values or distribution of A-B hopping terms, Peierls phases, and loss rates. This indicates that the bulk dynamics can be effectively captured in these models by specific spectrums characteristics under PBCs.
@@ IV.1 A-B coupling model 2-nd nearest intercell hopping
Figure 4: (a), (b) Energy spectrums under (a) PBCs and (b) OBCs, $t_{0}=0.3$ (red), $0.4$ (blue), and $0.5$ (green). (c), (d) The bulk distribution of $P_{x}$ in double logarithmic (c) and logarithmic (d) plots. System size $L=200$, and the initial position $x_{0}=150$. Through (a)-(d), $t_{p}=0.5$, $t_{1}=0.5$, $t_{2}=0.1$, $\gamma=0.5$ are fixed.
Figure 5: (a) The relative height $P_{\mathrm{edge}}/P_{\mathrm{min}}$ with varying $t_{2}$. (b) Part of the energy spectrums with system size $L=500$ under PBCs, $t_{2}=0.25$ (red), $0.33$ (blue), and $0.50$ (green) (c), (d) The loss probability $P_{x}$ for a walker initiated at $x_{0}=150$ with $L=200$. $t_{2}=0.2$ for (c) and $t_{2}=0.5$ for (d). Through (a)-(d), $t_{p}=0.5$, $t_{0}=0.3$, $t_{1}=0.5$, $\gamma=0.5$ are fixed.
@@
This is a quartic equation of $\beta$ with the constraint $|\beta|=1$. Therefore, the maximum number of IGC points is four, which is confirmed by the numerical results of the energy spectrums. Furthermore, we can infer that if intercell A-B couplings are up to the $n$-th nearest neighbors, the maximum number of IGC points is $2n$.
@@
with $\cos k_{0}=-t_{1}/4t_{2}$. Combining Eq. (23) and Eq. (23), we obtain the maximum value of $h_{x}(k_{0})$ as $t_{0}-(t_{1}/\sqrt{2})$, achieved when $t_{2}=t_{1}/2\sqrt{2}$. Thus, the inequality $h_{x}(k_{0})\leq 0$ always holds in the regime $t_{0}\leq t_{1}/\sqrt{2}$, regardless of the value of $t_{2}$. Consequently, the energy spectrums of the Bloch Hamiltonian Eq. (21) are guaranteed to exhibit IGC points. However, for $t_{0}>t_{1}/\sqrt{2}$, the existence of IGC points becomes contingent on specific values of hopping parameters $t_{0}$, $t_{1}$, and $t_{2}$. For example, as shown in Fig. 4(a), when fixing other parameters and increasing the parameter $t_{0}$, two IGC points gradually close together. At a critical value, they merge into a single IGC point. Besides, no IGC points exist when $t_{0}$ exceeds this critical value. Apparently, the corresponding OBC spectrums do not have IGC points, which is shown in Fig. 4(b). As discussed in Sec. III, whether energy spectrums have IGC points can be linked to the quantum walk behavior. Fig. 4(c) and 4(d) illustrate this connection, where the bulk distribution of particle loss probabilities $P_{x}$ exhibits power-law behavior for PBC spectrums with IGC points and exponential behavior for spectrums without them.
@@ IV.2 Model with varying Peierls phases $\phi$
Figure 6: (a) Energy spectrums under PBCs. Peierls phases $\phi$ are set to 00 (red), $\pi/6$ (black), $\pi/3$ (green), and $\pi/2$ (blue). (b) The relative height $P_{\mathrm{edge}}/P_{\mathrm{min}}$ with varying $\phi$. System size $L=200$ and initial position $x_{0}=150$. Throughout (a)-(b), $t_{p}=0.5$, $t_{0}=0.3$, $t_{1}=0.5$, $\gamma=0.5$ are fixed.
@@
This analytical expression exhibits good agreement with numerical results of energy spectrums under PBCs, as shown in Fig. 6(a). Therefore, Peierls phases $\phi$ do not alter the IGC conditions or impact the bulk scaling behavior of $P_{x}$. they only shift the values of $E_{IGC}$ along the real axis.
When $\phi=0$, the Hermitian part of the Bloch Hamiltonian Eq. (25) has time-reversal symmetry. Consequently, the onsite dissipation does not induce the NHSE. In this case, the distribution of $P_{x}$ is left-right symmetric, and there is no edge burst phenomenon. As $\phi$ increases from zero, this model breaks time-reversal symmetry and thus features the NHSE. We can see from Fig. 6(b) that the quantum walk exhibits a bipolar edge burst for a not very large $\phi$ and reduces to a standard edge burst on the left side when it exceeds some critical value. The corresponding PBC spectrums also exhibit the emergence and disappearance of self-intersecting points, as shown in Fig. 6(a).

### v2

# Imaginary gap-closed points and dynamics in a class of dissipative systems

@@ I Introduction
On the other hand, the NHSE also induces novel dynamical phenomena in open quantum systems. These include damping behavior and diverging relaxation time in Liouvillian dynamics [100, 101, 102, 103, 104], directional amplification of signals [105, 106], self-healing of skin modes [107], and directional invisibility of scattered wave-packet [108, 109]. Intriguingly, non-Hermitian dynamics exhibit boundary condition independence in the thermodynamic limit [110]. However, in finite-size systems, energy spectra under open boundary conditions (OBCs) show dramatic differences from those under periodic boundary conditions (PBCs) due to the NHSE [69, 97, 98]. Thus, one may question which non-Hermitian Hamiltonians determine the dynamical evolution of the systems. Furthermore, the boundary’s role is also crucial to understanding non-Hermitian dynamics.
Motivated by the above questions, we first analyze a general non-Hermitian model whose non-Hermiticity comes from the onsite dissipation. Our investigation centers on imaginary gap-closed (IGC) points, which are eigenstates possessing real energies and dictate the long-time behavior of the system [111]. Notably, these IGC points depend solely on the Hermitian component of the Hamiltonian, unaffected by dissipative terms. Furthermore, IGC states are populated on non-dissipative sites and can be viewed as dark modes. When focusing on a class of one-dimensional dissipative models, the PBC spectra exhibit IGC points, while their OBC counterparts lack such points as eigenstates feature the NHSE. By introducing the non-Hermitian quantum walk, we investigate the relation between IGC points and the scaling behavior of bulk loss probability in space. Consequently, these IGC points lead to a spatial power-law decay of the loss probability in the bulk of the system. Moreover, an associated boundary phenomenon, termed ”edge burst,” emerges. Specifically, when a quantum particle is initially prepared on a specific bulk site, there is a prominent peak in the loss probability at the edge. We discuss this phenomenon in different quantum walk scenarios and the impact of various parameters, showing the universality of the correlation between edge burst and IGC points. Finally, we analyze the Liouvillian dynamics and demonstrate that damping matrices of these dissipative chains under PBCs can have Liouvillian gapless points, implying an algebraic convergence towards the steady state in long-time dynamics.
@@ II Non-Hermitian Model and universal imaginary gapless modes
Let us first consider a general non-Hermitian lattice model [Fig. 1], whose non-Hermiticity totally comes from the onsite dissipation. We classify the lattice sites into two categories: non-dissipative sites and dissipative sites, denoted by $(h,i)$ and $(nh,i)$, respectively. Then, the Hamiltonian of this system can be divided into three parts,
@@
where $H_{h}$ is the interactions within non-dissipative sites, $H_{nh}$ is the interactions within dissipative sites, and $H_{inter}$ is the interactions between non-dissipative sites and dissipative sites. Their expressions are given by
@@
$H_{inter}=$ $\sum_{i,j}\left(c_{ij}\ket{nh,i}\bra{h,j}+h.c.\right),$
where coefficients $a_{ij}$, $b_{ij}$ and $c_{ij}$ are complex number, satisfying $a_{ij}=a_{ji}^{*}$, $b_{ij}=b_{ji}^{*}(i\neq j)$ and $\mathrm{Im}(b_{ii})<0$ for all $i$ and $j$. Thus, $H_{h}$ and $H_{inter}$ are Hermitian operators, while $H_{nh}$ is non-Hermitian.
@@
In systems where non-Hermiticity arises solely from onsite dissipation [Eq. (2)], the wave functions of IGC states are only populated on non-dissipative sites, i.e.
@@
where $-\gamma_{j}=\mathrm{Im}(b_{jj})$ is the dissipative rate of site $(nh,j)$, and $\gamma_{j}>0$. Notice that the LHS of Eq. (6) also equals to
@@
$(H_{h}-E_{IGC})\ket{IGC}=0,\quad H_{inter}\ket{IGC}=0.$ (9)
This implies that IGC states not only belong to the kernel of $H_{inter}$ but also function as eigenstates of a system solely comprised of non-dissipative sites, retaining their original eigenvalues. The term $H_{inter}$ can be interpreted as a representation of the coupling between non-dissipative and dissipative sites. Eq. (9) signifies that upon connecting the non-dissipative subsystem to the dissipative subsystem, these IGC states remain eigenstates of the combined system with their eigenvalues preserved.
In summary, IGC states can be described as dark modes characterized by the absence of a population on dissipative sites. Notably, IGC states are the eigenstates of $H_{h}$, satisfying the connection conditions $H_{inter}$, independent of $H_{nh}$. Our findings hold true for a broad class of systems, as evidenced by their applicability to arbitrary Hamiltonian that only have non-Hermitian terms on its diagonal like Eq. (2). However, limitations arise when considering effective Hamiltonians with off-diagonal non-Hermitian terms. These non-diagonal forms, often encountered in systems like atoms coupled by the electromagnetic field [112, 113, 114], whose dissipative dynamics can lead to phenomena like fluorescence [115], superradiance [116, 117, 118], and subradiance [119]. Our current theory necessitates further development to encompass these phenomena effectively.
@@ III Imaginary Gapless modes and dynamics of Quantum walk
where $k=2\pi n/L$ with $n=0,1,2,\cdots,L-1.$
The specific value of $k$ is determined by the connection condition obtained in Eq. (9)
@@
At $k=0$, the LHS of Eq. (12), denoted as $F(k)$, attains its maximum value $\sum_{m=0}^{n}t_{m}>0$. So, if the minimum value of $F(k)$ is smaller than or equal to zero, Eq. (12) has real roots, and the energy spectrum is IGC. In finite chains, satisfying condition (12) with a discrete crystal momentum $k$ might not be achievable. However, from a dynamical perspective, their bulk dynamical behavior mirrors that of the infinite-chain limit due to the local nature of propagation. Treating $k$ as continuous in this limit enables the identification of rigorous IGC points.
@@
Notably, the system under OBC lacks IGC states. This is because any IGC state would also be an eigenstate of the Hermitian $H_{h}$, which is an extended state, contradicting the NHSE’s localized nature under OBC. Alternatively, in systems with the NHSE, OBC spectra are enclosed by their PBC counterparts [98, 109]. For this dissipative lattice, the PBC spectra’ maximum imaginary part is zero, rendering the OBC spectra always imaginary-gapped. An example is shown in Fig. 3(a).
Although substantial differences exist in the spectrum and eigenstates between OBC and PBC for this model, the dynamics within the bulk exhibit similarity for both cases. This similarity is evident from the Schrödinger equation $i(d/dt)\ket{\psi(t)}=H\ket{\psi(t)}$ (we set $\hbar=1$ throughout this paper). Under PBC, the Hamiltonian is given by $H_{PBC}=H_{OBC}+\delta H$. Here, $\delta H$ represents the coupling term that connects the two boundaries. When a wave function |$\ket{\psi(t)}$ predominantly resides in the bulk, with negligible population at the boundaries, $\delta H\ket{\psi(t)}\approx 0$, and then consequently, the Schrödinger equations for OBC and PBC become effectively equivalent in this approximation. This feature is exemplified in Fig. 3(b). In the following, we show that the dynamical evolution properties of quantum walks under OBCs can be elucidated by IGC points of the PBC spectra.
Figure 3: The model with only the nearest A-B coupling (a) Energy spectra with $t_{0}=0.3$. Blue: OBC; Red: PBC. (b) The bulk distribution of the relative loss probabilities $P_{x}/P_{130}$ in a double logarithmic plot with $t_{0}=0.3$ for system size $L=200$, initial position $x_{0}=150$. (c), (d) The distribution of $P_{x}$ for a walker initiated at $x_{0}=150$ with $L=200$ under OBC. $t_{0}=0.3$ for (c) and $t_{0}=0.6$ for (d). (e) Relative height $P_{\rm{edge}}/P_{\min}$ and (f) the edge loss probability $P_{\rm{edge}}$ with chain length $L=200$, and $x_{0}$ varying from 40 to 160. $t_{0}=0.3$ for (c) and $t_{0}=0.6$ for (d) in a logarithmic plot. Throughout (a)-(f), $\phi=\pi/2$, $t_{1}=0.5$, $t_{p}=0.5$, and $\gamma=0.5$ are fixed.
@@
Note that $\Sigma_{x}P_{x}=1$ is satisfied under the initial-state normalization $\left\langle\psi(0)|\psi(0)\right\rangle=1$. Intuitively, the distribution $P_{x}$ features a peak centered on the site $x_{0}$, displaying left-right asymmetry attributable to the NHSE, which causes the wave function to move leftward preferentially [123]. However, under certain parameters, a prominent peak, named edge burst, emerges at the left edge [124]. In a recent paper, Xue et al. investigated this model with the nearest A-B hopping term and uniform dissipation, i.e., $\gamma_{x}$ is a constant, revealing that the edge burst phenomenon results from the interplay between the IGC spectrum under PBC and the NHSE [111].
@@ IV Impact of parameters
The above discussion shows that IGC points of PBC spectra are crucial to the non-Hermitian dynamics. In this section, we explore the quantum walk across various parameters, demonstrating the universality of the correlation between edge burst and IGC points. This correlation remains independent of the values or distribution of A-B hopping terms, Peierls phases, and loss rates. This indicates that the bulk dynamics can be effectively captured in these models by specific spectra characteristics under PBCs.
@@ IV.1 A-B coupling model 2-nd nearest intercell hopping
Figure 4: (a), (b) Energy spectra under (a) PBCs and (b) OBCs, $t_{0}=0.3$ (red), $0.4$ (blue), and $0.5$ (green). (c), (d) The bulk distribution of $P_{x}$ in double logarithmic (c) and logarithmic (d) plots. System size $L=200$, and the initial position $x_{0}=150$. Through (a)-(d), $t_{p}=0.5$, $t_{1}=0.5$, $t_{2}=0.1$, $\gamma=0.5$ are fixed.
Figure 5: (a) The relative height $P_{\mathrm{edge}}/P_{\mathrm{min}}$ with varying $t_{2}$. (b) Part of the energy spectra with system size $L=500$ under PBCs, $t_{2}=0.25$ (red), $0.33$ (blue), and $0.50$ (green) (c), (d) The loss probability $P_{x}$ for a walker initiated at $x_{0}=150$ with $L=200$. $t_{2}=0.2$ for (c) and $t_{2}=0.5$ for (d). Through (a)-(d), $t_{p}=0.5$, $t_{0}=0.3$, $t_{1}=0.5$, $\gamma=0.5$ are fixed.
@@
This is a quartic equation of $\beta$ with the constraint $|\beta|=1$. Therefore, the maximum number of IGC points is four, which is confirmed by the numerical results of the energy spectra. Furthermore, we can infer that if intercell A-B couplings are up to the $n$-th nearest neighbors, the maximum number of IGC points is $2n$.
@@
with $\cos k_{0}=-t_{1}/4t_{2}$. Combining Eq. (23) and Eq. (24), we obtain the maximum value of $h_{x}(k_{0})$ as $t_{0}-(t_{1}/\sqrt{2})$, achieved when $t_{2}=t_{1}/2\sqrt{2}$. Thus, the inequality $h_{x}(k_{0})\leq 0$ always holds in the regime $t_{0}\leq t_{1}/\sqrt{2}$, regardless of the value of $t_{2}$. Consequently, the energy spectra of the Bloch Hamiltonian Eq. (21) are guaranteed to exhibit IGC points. However, for $t_{0}>t_{1}/\sqrt{2}$, the existence of IGC points becomes contingent on specific values of hopping parameters $t_{0}$, $t_{1}$, and $t_{2}$. For example, as shown in Fig. 4(a), when fixing other parameters and increasing the parameter $t_{0}$, two IGC points gradually close together. At a critical value, they merge into a single IGC point. Besides, no IGC points exist when $t_{0}$ exceeds this critical value. Apparently, the corresponding OBC spectra do not have IGC points, which is shown in Fig. 4(b). As discussed in Sec. III, whether energy spectra have IGC points can be linked to the quantum walk behavior. Fig. 4(c) and 4(d) illustrate this connection, where the bulk distribution of particle loss probabilities $P_{x}$ exhibits power-law behavior for PBC spectra with IGC points and exponential behavior for spectra without them.
@@ IV.2 Model with varying Peierls phases $\phi$
Figure 6: (a) Energy spectra under PBCs. Peierls phases $\phi$ are set to 00 (red), $\pi/6$ (black), $\pi/3$ (green), and $\pi/2$ (blue). (b) The relative height $P_{\mathrm{edge}}/P_{\mathrm{min}}$ with varying $\phi$. System size $L=200$ and initial position $x_{0}=150$. Throughout (a)-(b), $t_{p}=0.5$, $t_{0}=0.3$, $t_{1}=0.5$, $\gamma=0.5$ are fixed.
@@
This analytical expression exhibits good agreement with numerical results of energy spectra under PBCs, as shown in Fig. 6(a). Therefore, Peierls phases $\phi$ do not alter the IGC conditions or impact the bulk scaling behavior of $P_{x}$. they only shift the values of $E_{IGC}$ along the real axis.
When $\phi=0$, the Hermitian part of the Bloch Hamiltonian Eq. (25) has time-reversal symmetry. Consequently, the onsite dissipation does not induce the NHSE. In this case, the distribution of $P_{x}$ is left-right symmetric, and there is no edge burst phenomenon. As $\phi$ increases from zero, this model breaks time-reversal symmetry and thus features the NHSE. We can see from Fig. 6(b) that the quantum walk exhibits a bipolar edge burst for a not very large $\phi$ and reduces to a standard edge burst on the left side when it exceeds some critical value. The corresponding PBC spectra also exhibit the emergence and disappearance of self-intersecting points, as shown in Fig. 6(a).